OpenStudy (anonymous):

Write an equation of a line that passes through (9,-5) and (3,-5).

OpenStudy (solomonzelman):

first find the slope of this line, do you know how to do that?

OpenStudy (solomonzelman):

\[\frac{y_1-y_2}{x_1-x_2}=slope\]

OpenStudy (solomonzelman):

plug in your points into (the above) slope formula.

OpenStudy (jdoe0001):

\(\begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(9\quad ,&-5)\quad &(3\quad ,&-5) \end{array} \\\quad \\ \large slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ \bf y-y_1=m(x-x_1)\quad \textit{plug in your values to get the equation}\)

OpenStudy (solomonzelman):

@cuitiefromthesouth2, we need your cooperation to help you, we can't just do your work, please say something.....

OpenStudy (anonymous):

sorry I was trying to work out the problem

OpenStudy (solomonzelman):

Oh, go a head and work in here PLEASE! We will correct you if your wrong.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

m=-5-(-5)/3-9

OpenStudy (anonymous):

=0/6

OpenStudy (anonymous):

0/-6

OpenStudy (solomonzelman):

cuitiefromthesouth2, the other way, not -5-(-5)/3-9, IT IS 3-9/-5-(-5)

OpenStudy (anonymous):

=-6/0

OpenStudy (solomonzelman):

yep! The slope is undefined and REMEMBER: if slope is undefined the line is vertical, because in (any) vertical the x coordinates are all same. lets take a vertical line

OpenStudy (solomonzelman):

See?

OpenStudy (anonymous):

yea , thank you

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

So how would you write the equation of your line?

OpenStudy (solomonzelman):

In my example the equation would be x=a, what about your's?

OpenStudy (anonymous):

3-9/-5-(-5) = -6/0, which is undefined ?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

yes, so it is x=-6